OpenStudy (anonymous):

find the x-coordinante of the vertex. x^2+12x+21

OpenStudy (anonymous):

the formula h = -b / 2a to find the x coordinate h of the vertex of a parabola and k = f(h) to find the y coordinate k of the vertex. To find the x intercepts, it solves the quadratic equation ax2 + bx + c = 0. The y intercept if given by f(0). so, x coordinate h of the vertex, h = -b / 2a, h=-12/2=-6